OpenStudy (anonymous):

A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. What is the original number?

OpenStudy (anonymous):

x-7(3)=-9 /3 /3 x-7=-3 +7 +7 x=4

OpenStudy (anonymous):

thanks! @jprose

OpenStudy (anonymous):

np